OpenStudy (anonymous):

This graph of a function is a translation of y =4/x . What is an equation for the function? Can someone help me pleasee?

OpenStudy (zehanz):

Remember: if you translate y = f(x) by a horizontal units and b vertical units, you get y - b = f(x-a), assuming right and up are positive values. So if you have y = x² and translate it 3 to the right and 4 up, you get: y - 4 = (x-3)², so the formula for the translated graph is y = (x-3)² + 4. Now try this method with y=4/x!

OpenStudy (anonymous):

umm im still kind of confused

OpenStudy (zehanz):

I see why now. y=4/x has already been translated; from "this function". Shouldn't you have to know the graph of "this function"?

OpenStudy (anonymous):

yeah i can see the graph but im still confused do you wanna see it?

OpenStudy (zehanz):

Yes please

OpenStudy (zehanz):

If you start with \[f(x)=\frac{ 4 }{ x }\]then translate the graph 6 units to the left (a = -6) and two up (b = 2), you get according to the formula above: g(x) - 2 = f(x - -6), so g(x) = f(x+6)+2. With the function 4/x this is:\[g(x)=\frac{ 4 }{ x+6 }+2\]See attached image!

OpenStudy (anonymous):

ohh okay i get it now thank you soo much!!! :)